Charge 1 is positive and is located at the origin and charge 2 is negative. The electric force on charge 1 due to charge 2 is 292.2 degrees measured counter-clockwise from the +x axis. What is the direction of charge 2 due to charge 1 measured counter-clockwise from the +x axis, in degrees? Nothing related to the charges, changes.... they are still the same charges in the same positions.
Answer #1

Solution :- Given data:-

The electric force on charge 1 due to charge 2 is 292.2 degrees

From Newton's third law,

For every action there is equal and opposite reaction.

Force on charge 2 due to charge 1 = -Force on charge 1 due to charge 1

here -ve sign represents that direction of force will be opposite.

Now given that

Direction of Force on charge 1 due to charge 2 = 292.2 °CCW from +x-axis

Direction of Force on charge 2 due to charge 1 = \theta deg CCW from +x-axis

Now these forces are in opposite directions, So

292.2- \theta = 180 deg

\theta = 292.2- 180

\theta = 112.2 deg Counter-clockwise from +x axis